# SubTrack your Grad: Gradient Subspace Tracking for Memory and Time Efficient Full-Parameter LLM Training

### v1

@@ Abstract
Training Large Language Models (LLMs) demand significant time and computational resources due to their large model sizes and optimizer states. To overcome these challenges, recent methods, such as BAdam, employ partial weight updates to enhance time and memory efficiency, though sometimes at the cost of performance. Others, like GaLore, focus on maintaining performance while optimizing memory usage through full parameter training, but may incur higher time complexity. By leveraging the low-rank structure of the gradient and the Grassmannian geometry, we propose SubTrack-Grad, a subspace tracking-based optimization method that efficiently tracks the evolving gradient subspace by incorporating estimation errors and previously identified subspaces. SubTrack-Grad delivers better or on-par results compared to GaLore, while significantly outperforming BAdam, which, despite being time-efficient, compromises performance. SubTrack-Grad reduces wall-time by up to 20.57% on GLUE tasks (15% average reduction) and up to 65% on SuperGLUE tasks (22% average reduction) compared to GaLore. Notably, for a 3B parameter model, GaLore incurred a substantial 157% increase in wall-time compared to full-rank training, whereas SubTrack-Grad exhibited a 31% increase—representing a 49% reduction in wall-time, while enjoying the same memory reductions as GaLore.
Machine Learning, ICML
Figure 1: Visualization of SubTrack-Grad: Between subspace updates, gradients are projected onto a fixed subspace. The tangent vector $\nabla F$ is computed via the derivative of a loss function that measures the subspace estimation error. The subspace is then updated by moving along the corresponding geodesic, determined by $\nabla F$, on the Grassmannian manifold to minimize estimation error.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated state-of-the-art performance across a wide range of tasks and are rapidly growing in popularity. However, training and fine-tuning these models require significant resources, including extensive hardware and time, which limits their practicality for many applications and increases their environmental impact due to a larger carbon footprint. (Zhao et al., 2024; Jaiswal et al., 2024; Muhamed et al., 2024; Miles et al., 2024; Modoranu et al., 2024; Hao et al., 2024; Li et al., 2024). Consequently, there is a pressing need to develop time- and memory-efficient methods to make LLMs more accessible and reduce their environmental impact, while preseving performance.
Several techniques have been proposed to address memory challenges, including gradient checkpointing (Chen et al., 2016) and memory offloading (Rajbhandari et al., 2020). In this context, several approaches aim to reduce memory usage by optimizing a subset of model parameters or operating within a lower-dimensional space (Dettmers et al., 2024; Yaras et al., 2024; Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Miles et al., 2024; Hu et al., 2021). Notably, the well-known method LoRA (Hu et al., 2021), addresses memory efficiency by decomposing weight matrices into two low-rank trainable matrices, enabling the optimization of network parameters within a smaller subspace and significantly reducing memory usage. However, LoRA’s effectiveness relies on the assumption that parameters reside in a low-dimensional space, which does not always hold true. On the other hand, recent methods including BAdam (Luo et al., 2024) and the one introduced by Ramesh et al. (2024) leverage the block coordinate descent (BCD) framework and partition the model into different blocks to optimize one block at a time. These methods can effectively reduce memory footprint while avoiding extra computation; however, while these achieve better memory and time efficiency, they suffer from performance degradation due to partial parameter tuning.
Memory requirements extend beyond trainable parameters, with a significant portion consumed by optimizers for storing element-wise states (Zhao et al., 2024). To address this, recent efforts have focused on reducing the optimizer parameters while still targeting full parameter training (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Zhang et al., 2024; Modoranu et al., 2024; Zhao et al., 2024; Muhamed et al., 2024). Prominantly, GaLore (Zhao et al., 2024) reduces the memory consumption of optimizers by projecting gradient matrices into a low-rank subspace and tracking changes through periodic singular value decomposition (SVD) to obtain a rank-$r$ approximation. Unlike weight parameters, there are many works showing that gradients evolve in a low-dimensional space during gradient descent (Gur-Ari et al., 2018; Schneider et al., 2024; Yaras et al., 2023); using this fact, GaLore decreases optimizer memory footprint while enabling full and simultaneous parameter updates. However, their approach faces several challenges. First, SVD is computationally expensive, and if the gradient does not evolve within a nearly constant subspace, GaLore must increase the frequency of SVD operations, significantly increasing the amount of computation. This is problematic because not all layers’ gradients converge to a stable subspace early in training (Jaiswal et al., 2024). Moreover, applying SVD to a single gradient matrix is susceptible to data noise (Vaswani et al., 2018), and GaLore does not leverage 1) information from the orthogonal space as feedback to adjust the subspace (Modoranu et al., 2024) or 2) previously computed subspaces to incorporate past knowledge, which could help mitigate noise effects and improve convergence.
To this end, we propose Gradient Subspace Tracking (SubTrack-Grad), a Grassmannian-based, time- and memory-efficient method that adjusts the subspace using rank-$1$ updates for full parameter fine-tuning. SubTrack-Grad leverages information from the orthogonal complement of the subspace to enhance subspace estimation through simple linear operations, making it more computationally efficient than GaLore by avoiding periodic SVD on gradient matrices while maintaining the same memory efficiency. Furthermore, its ability to perform simultaneous weight updates provides significant performance improvements over BAdam. Finally, SubTrack-Grad dynamically adapts to the gradient subspace changes, reducing abrupt shifts for faster convergence, in some cases achieving better performance than full-rank finetuning! An overview of SubTrack-Grad is provided in Figure 1, and our main contributions are as follows:
  • We introduce SubTrack-Grad, a memory- and time-efficient full-parameter training that maintains performance by tracking the gradient subspace using Grassmannian geometry.
  • We show that SubTrack-Grad outperforms BAdam, and delivers performance on par with, or superior to GaLore, while significantly reducing runtime with maintaining GaLore’s memory gains.
  • We show that tracking the gradient subspace helps reduce abrupt changes in the optimization process, thereby accelerating convergence.
  • We prove that our method aligns with GaLore’s convergence guarantees while enabling more frequent subspace updates by exercising greater control over subspace adjustments through the use of prior knowledge and estimation errors.
2 Related Works
Several works aim to improve the efficiency of training and fine-tuning LLMs, addressing a growing demand as their popularity rapidly increases. LoRA (Hu et al., 2021), a widely recognized method for reducing the number of trainable parameters, that projects model weights into a lower-dimensional space, resulting in two trainable low-rank matrices. This approach significantly reduces memory requirements for fine-tuning LLMs. Dettmers et al. (2024) build on LoRA by employing quantization techniques and paged optimizers to further reduce memory usage. Additionally, Yaras et al. (2024) introduce Deep LoRA, which uses deep matrix factorization for low-rank optimization, addressing overfitting issues and reducing the need for precise tuning of the rank parameter. Several other works have also extended LoRA to enhance the efficiency of training and fine-tuning LLMs (Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Pan et al., 2024). Miles et al. (2024) propose compressing intermediate activation vectors and reconstructing them during backpropagation to enhance memory efficiency. Additionally, Hao et al. (2024) demonstrate that full-parameter fine-tuning is feasible by random projections on the gradient matrix, showing that LoRA essentially performs a down-projection of the gradient. BAdam (Luo et al., 2024) leverages the block coordinate descent (BCD) framework to reduce memory consumption while maintaining capabilities comparable to Adam.
Several approaches aim to reduce memory consumption in optimizers, as optimizers like Adam (Kingma & Ba, 2017) account for a significant portion of memory usage due to their storage of element-wise states to improve the optimization process (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Rajabi & Rambhatla, 2024c, a). MicroAdam (Modoranu et al., 2024) tackles this issue by compressing the gradient space for optimization and utilizing the resulting compression error through feedback loops to improve the optimization process. Adam-mini (Zhang et al., 2024) partitions model parameters into blocks, assigning a single learning rate to each block. This design achieves a significant reduction in memory usage while maintaining performance. Gur-Ari et al. (2018) suggests that a substantial portion of gradients lies within a small, largely consistent subspace, a finding also reported by other studies, including Schneider et al. (2024); Yaras et al. (2023). GaLore (Zhao et al., 2024) leverages this property to reduce the memory requirements of optimizers by projecting gradients into a lower-dimensional subspace and then projecting them back for full parameter tuning. This approach has been integrated with other methods to further reduce memory usage during training and fine-tuning of LLMs (Li et al., 2024; Rajabi & Rambhatla, 2024b). However, not all layers’ gradients evolve within a stable low-rank subspace. Jaiswal et al. (2024) identifies layers with constantly changing gradients where low-rank projection may be inefficient for tuning. By analyzing the distribution of singular values across different layers, they select those that evolve within a small subspace for fine-tuning while freezing the remaining layers. Gradient Structured Sparsification (Grass) (Muhamed et al., 2024) further reduces memory usage by applying sparse projection matrices to the gradient, transforming the gradient matrix into a sparse space. This approach leverages sparse representations to significantly decrease the memory footprint. In Ramesh et al. (2024) the authors propose an approach that dynamically selects and updates a small subset of parameters, for a faster and more memory-efficient training without altering the model’s structure.
A common approach in working with high-dimensional data is to project the data into a lower-dimensional space, and many studies focus on tracking these subspaces as they evolve. Balzano et al. (2011) introduce an incremental method for updating subspaces on the Grassmannian when the data is partially observed. Zhang & Balzano (2016) and Kasai (2017) propose methods to handle noise effect in tracking these subspaces. Furthermore, Blocker et al. (2023) present a method for evolving geodesic-based data in the Grassmannian for updating the subspace effectively. Also, Rajabi & Rambhatla (2024d) have applied Grassmannian-based subspace tracking on gradient matrices to increase the efficiency of training.
3 SubTrack-Grad: Tracking the Gradient Subspace
Since gradients typically evolve within a small subspace, compressing this space can significantly reduce the memory footprint of the optimizer. As demonstrated in Zhao et al. (2024) whenever the gradient takes the general form
$\small G=\sum_{i}A_{i}+\sum_{i}B_{i}WC_{i}$ (1)
where $i$ denotes the batch index, and $B_{i}$ and $C_{i}$ are positive semi-definite (PSD) matrices, this gradient can be projected onto a small subspace that remains nearly stable while ensuring that the optimization process continues to converge, as discussed in Section 4 and Appendix A. However, the subspace of the gradient is not always stable, making it crucial to track its changes. GaLore (Zhao et al., 2024) addresses this by periodically performing SVD on gradient matrices, while keeping the update frequency low to align with the assumption of a stable subspace, which is used in its convergence guarantee. This approach poses several challenges: 1) not all gradients converge to a stable subspace within a few iterations, 2) SVD is computationally expensive, and increasing the frequency of updates to capture changes significantly raises both run-time and environmental costs, and 3) increasing the update frequency contradicts the assumption of a stable subspace, as SVD is sensitive to noise and does not account for the previously computed subspace or estimation error to regulate the extent of change applied.
We propose SubTrack-Grad, a memory- and time-efficient method that allows full and simultaneous parameter tuning. SubTrack-Grad utilizes the estimation error and the previously computed subspace to adjust the core subspace. It effectively controls the amount change, allowing more frequent updates without compromising the stability assumption; while, as illustrated in Figure 4, increasing the update frequency will not drastically increase the runtime. The subspace is initialized using SVD as follows:
$\small\begin{split}G_{0}=USV^{\top}\approx\sum_{i=1}^{r}s_{i}u_{i}v_{i}^{\top},\\ P_{0}=[u_{1},u_{2},...,u_{r}],\quad Q_{0}=[v_{1},v_{2},...,v_{r}].\end{split}$ (2)
Here, $G_{0}$ is an $m\times n$ gradient matrix at step 00, $U$, $S$, and $V$ are its SVD components, and $r$ is the specified rank. In each optimization step, the gradients are projected onto the subspace of the left singular vectors if $m\leq n$, or onto the right singular vectors otherwise, thereby optimizing memory usage (Zhao et al., 2024). After the optimizer processes the low-rank gradient, it gets projected back for a normal weight update. From now on, we assume $m\leq n$ without loss of generality, implying that $S_{0}=P_{0}$, an $m\times r$ orthonormal matrix whose columns span the underlying subspace.
At each step $t$, the matrix $S_{t}$ represents the subspace at that step, and projects the gradient matrix $G_{t}$ onto that subspace by $\widetilde{G}_{t}=S_{t}^{\top}G_{t}$; where $\widetilde{G}_{t}$, the projection of the gradient onto a rank-$r$ subspace, will be an $r\times n$ matrix. The optimizer then performs within this low-rank space, which substantially reduces the number of state parameters and thus the memory footprint. The optimizer outputs $\widetilde{G}_{t}^{O}$, which is then projected back to the original space by $\widehat{G}_{t}=S_{t}\rho_{t}\widetilde{G}_{t}^{O}$, where $\rho_{t}$ represents the entry-wise regularizer used in the optimizer, to be passed to the network.
Algorithm 1 SubTrack-Grad
0:  Sequence of $m\times n$ gradients $G_{t}$ with $m\leq n$ (w.l.o.g.), step-size $\eta$, rank $r$, subspace update steps k
  Initialize Subspace via SVD Decomposition: $P_{0}\leftarrow U[:,:r]$ , where $U,S,V\leftarrow\text{SVD}(G_{0})$ $S_{0}\leftarrow P_{0}$
  for $t=1,\ldots,T$ do
     if $t\mod k=0$ then
       Update subspace: $G_{lr}=\arg\min_{A}\|(S_{t-1}A-G_{t})\|^{2}$ $R=G_{t}-S_{t-1}G_{lr}$ $\nabla F=-2RG_{lr}^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ $S_{t}=(S_{t-1}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t-1}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
     else
       Keep using previous subspace: $S_{t}=S_{t-1}$
     end if
     Return final projected gradient to the optimizer: $S_{t}^{\top}G_{t}$
  end for
As previously discussed, the gradient does not always evolve within a stable subspace; hence, $S_{t}$, the orthonormal matrix spanning the core subspace, must be appropriately updated. SubTrack-Grad updates this subspace by moving along Grassmannian geodesics, and leverages the previously computed subspace and the current estimation error to minimize abrupt changes and noise effects, as illustrated in Figure 3.
We frame the problem of identifying the subspace as selecting the appropriate element from the Grassmannian, the set of all $d$-dimensional subspaces within an $n$-dimensional vector space (Bendokat et al., 2024). The subspace is updated after $k$ steps, the subspace update interval, and our objective is to minimize the Euclidean distance between the subspace and the current observed gradient in an update step. The corresponding cost function for this aim is defined as:
@@ 3 SubTrack-Grad: Tracking the Gradient Subspace
where $A$ is the solution to the least squares problem. The derivative of (3) with respect to $S_{t}$ is given in (4), and the residual $R={G_{t}}-S_{t}A$ lies in the orthogonal complement of $S_{t}$. To update the subspace in the appropriate direction, we calculate the tangent vector $\nabla F$ on the Grassmannian, as shown in (5) based on Edelman et al. (1998), where the second equality holds because $R$ is orthogonal to $S_{t}S_{t}^{\top}$.
@@
The tangent vector $\nabla F$ provides the direction for adjusting the subspace by accounting for the error lying in the orthogonal complement. However, to minimize changes to the subspace, SubTrack-Grad first computes a rank-$1$ approximation of $\nabla F$, determined by its largest singular value and the corresponding singular vector obtained from its SVD, and represented as $\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ in the final equality of (5). This approximation is then used to update the subspace.
@@
This update rule preserves the orthonormality of $S_{t+1}$, ensuring it remains on the Grassmannian. The last term in (5), is required when using thin-SVD; it projects the previous subspace onto the orthogonal complement of $\widehat{V}_{F}$, ensuring that the portion of $S_{t}$ which has not been updated in this step is still included. By leveraging the geometry of the Grassmannian, SubTrack-Grad effectively tracks the underlying subspace of the gradient space, and the pseudo-code for this method is provided in Algorithm 1.

### v2

# SubTrack++ : Gradient Subspace Tracking for Scalable LLM Training

@@ Abstract
Training large language models (LLMs) is highly resource-intensive due to their massive number of parameters and the overhead of optimizer states. While recent work has aimed to reduce memory consumption, such efforts often entail trade-offs among memory efficiency, training time, and model performance. Yet, true democratization of LLMs requires simultaneous progress across all three dimensions. To this end, we propose SubTrack++ that leverages Grassmannian gradient subspace tracking combined with projection-aware optimizers, enabling Adam’s internal statistics to adapt to changes in the optimization subspace. Additionally, employing recovery scaling, a technique that restores information lost through low-rank projections, further enhances model performance. Our method demonstrates SOTA convergence by exploiting Grassmannian geometry and achieves lowest evaluation loss, outperforming the current SOTA while reducing pre-training wall time by 43% and maintaining the memory footprint on a 1B-parameter Llama model.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated state-of-the-art performance across a wide range of tasks and are rapidly growing in popularity. However, training and fine-tuning these models require significant resources, including extensive hardware and time, which limits their practicality for many applications and increases their environmental impact and carbon footprint. (Zhao et al., 2024; Jaiswal et al., 2024; Muhamed et al., 2024; Miles et al., 2024; Modoranu et al., 2024; Hao et al., 2024; Li et al., 2024).
Several techniques have been proposed to mitigate memory bottlenecks, such as gradient checkpointing (Chen et al., 2016) and memory offloading (Rajbhandari et al., 2020). LoRA (Hu et al., 2021) and other low-rank adaptation methods (Dettmers et al., 2024; Hu et al., 2021; Yaras et al., 2024; Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Miles et al., 2024) have gained popularity by optimizing a reduced set of parameters. Such approaches often assume that the full parameter space is inherently low-rank—an assumption that does not always hold and can lead to suboptimal performance. In addition, methods like BAdam (Luo et al., 2024) and Block-LLM (Ramesh et al., 2024), utilize block coordinate descent to optimize parameter subsets, achieving memory savings at the cost of partial parameter tuning and reduced accuracy.
Figure 1: We compare baselines on pre-training a 1B-parameter model. (a) SubTrack++ achieves the lowest evaluation loss across all methods. (b) Its peak memory usage is significantly lower than APOLLO and LDAdam, and on par with GaLore and Fira. (c) In terms of wall-time, SubTrack++ incurs minimal overhead relative to APOLLO and is markedly faster than GaLore, Fira, and LDAdam. Overall, SubTrack++ outperforms all baselines in evaluation loss while matching or exceeding them in memory and runtime efficiency.
However, memory requirements extend beyond trainable parameters, with a significant portion consumed by optimizers for storing element-wise states (Zhao et al., 2024). To address this, recent efforts have focused on reducing the optimizer parameters while still targeting full parameter training (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Zhang et al., 2024; Modoranu et al., 2024; Zhao et al., 2024; Muhamed et al., 2024; Rajabi and Rambhatla, 2024a). Leveraging the fact that gradients evolve in a low-dimensional space during gradient descent (Gur-Ari et al., 2018; Schneider et al., 2024; Yaras et al., 2023) , GaLore (Zhao et al., 2024) reduces memory usage by projecting gradients into a low-rank subspace and periodically updating the rank-$r$ approximation via singular value decomposition (SVD)—an approach also used in more recent method, Fira (Chen et al., 2025). While SVD offers optimal low-rank approximation (Robert et al., 2025), performing it repeatedly can pose several challenges. First, it is computationally intensive, and alternatives like random projections (as in APOLLO (Zhu et al., 2025) and GoLore (He et al., 2025)) or approximation methods for estimating dominant singular values (as in LDAdam (Robert et al., 2025), Online Subspace Descent (Liang et al., 2024) and Rajabi and Rambhatla (2024b)) match or outperform SVD in practice. Moreover, SVD is sensitive to noise (Vaswani et al., 2018; He et al., 2025) and tends to degrade in late training stages when gradients are small, often hindering convergence (He et al., 2025).
Geometry-based methods have demonstrated strong performance in various machine learning applications (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Rajabi and Rambhatla, 2024c; Blocker et al., 2023). In this context, using Grassmannian manifolds for subspace tracking has led to structurally embedded information, lower computational complexity, and improved performance (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Blocker et al., 2023; Chakraborty et al., 2017; Zhang and Balzano, 2016). Grassmannian is the set of $r$-dimensional subspaces in a d-dimensional space, and the ultimate goal of low-rank optimization is to find the proper subspace in this space. This line of work has demonstrated robustness and efficiency in high-dimensional, noisy environments (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Balzano et al., 2018; Chakraborty et al., 2017), outperforming online or robust Principle Component Analysis (PCA) techniques (Li, 2004; Balzano et al., 2018; Chakraborty et al., 2017). Their natural robustness against perturbations and strong theoretical guarantees make them particularly well-suited to tracking the evolving gradient subspaces encountered in LLM training. (Balzano, 2012; Balzano et al., 2011).
Figure 2: Visualization of Grassmannian-based subspace tracking: Between subspace updates, gradients are projected onto a fixed subspace. The tangent vector $\nabla F$ is computed via the derivative of a loss function that measures the subspace estimation error. The subspace is then updated by moving along the corresponding geodesic, determined by $\nabla F$, on the Grassmannian manifold to minimize estimation error.
To this end, we employ subspace tracking on Grassmannian geodesics (Figure 2), to develop a geometry-based, time- and memory-efficient method for tracking the gradient subspace. This way, instead of reconstructing low-rank approximations via expensive SVD or random projections, we can efficiently leverage previously computed subspaces and the estimation error to better adjust the projection. We also incorporate subspace shifts into Adam’s first and second momentum update rules, ensuring proper alignment with coordinate changes using a Projection-Aware Optimizer (Robert et al., 2025). Additionally, we recover and scale (Recovery Scaling) the gradient information lost during low-rank projection to boost performance by utilizing the scaling information of the low-rank optimizer Chen et al. (2025); Zhu et al. (2025).
To summarize, SubTrack++ is a projection-aware geometry-based approach that supports full-parameter training and incorporates recovery scaling, offering superior time efficiency compared to SVD or PowerSGD methods (e.g., GaLore (Zhao et al., 2024), FiRA (Chen et al., 2025), LDAdam (Robert et al., 2025)), while maintaining GaLore’s memory footprint. It outperforms online PCA methods like Online Subspace Descent (Liang et al., 2024) in subspace tracking and achieves SOTA convergence and evaluation loss across all strong baselines; Comparison provided in Figure 1.
2 SubTrack++
Challenges of Low-Rank Optimization. Projecting gradients into a low-rank subspace reduces memory overhead and enables scalable LLM training, but it introduces important trade-offs. First, gradient subspaces can be unstable and require adaptive tracking; while SVD-based methods can capture these shifts (Zhao et al., 2024; Chen et al., 2025), they are computationally expensive. To address this, recent work has explored cheaper approximations and random projections (Robert et al., 2025; Liang et al., 2024; Zhu et al., 2025; He et al., 2025). Furthermore, optimizers like Adam assume a fixed coordinate system, so any changes in the subspace must be reflected in their internal states to preserve consistent momentum updates. Finally, low-rank projections inherently discard some gradient components, that recovering and utilizing these discarded signals can lead to improved performance (Chen et al., 2025; Zhu et al., 2025).
Overview. We propose SubTrack++ , a memory- and time-efficient method for full, simultaneous parameter tuning that embeds geometric insights into low-rank optimization. It improves efficiency and performance via three core components: 1) Subspace Tracking—Grassmannian-based updates that refine projections using estimation error and subspace history; 2) Projection-Aware Optimizer—adjusting Adam’s internal states to account for shifting subspaces; and 3) Recovery Scaling—rescaling discarded gradient components to reclaim lost information.
Subspace Tracking. In SubTrack++ , subspace estimation is framed as selecting a point on the Grassmannian—the manifold of all $d$-dimensional subspaces in an $n$-dimensional space (Bendokat et al., 2024)—which is the formal definition and goal of low-rank projections. This geometric perspective offers three key benefits: 1) By moving along Grassmann geodesics, it refines the subspace using prior subspace and estimation error, avoiding full reinitialization as done in (Zhao et al., 2024; Chen et al., 2025; He et al., 2025). 2) Updates rely on lightweight algebraic operations, minimizing time and memory costs. 3) Controlled subspace shifts improve robustness against noise and abrupt changes.
The initial subspace is computed using SVD as shown in (1). $G_{0}\in\mathbb{R}^{m\times n}$ is the gradient matrix at step 00; $U$, $S$, and $V$ are its SVD components, and $r$ is the specified rank.
$\small G_{0}=USV^{\top}\approx\sum_{i=1}^{r}s_{i}u_{i}v_{i}^{\top}$ (1)
At each step, gradients are projected onto the subspace of left singular vectors if $m\leq n$, and right singular vectors otherwise; optimizing memory usage (Zhao et al., 2024). We assume $m\leq n$ without loss of generality, so the subspace is represented by $S_{t}\in\mathbb{R}^{m\times r}$ ($S_{0}=[u_{1},...,u_{r}]$), an orthonormal basis spanning the top-$r$ directions. The gradient is projected as $\widetilde{G}_{t}=S_{t}^{\top}G_{t}\in\mathbb{R}^{r\times n}$, and the optimizer operates in this reduced space, significantly lowering memory and state overhead.
To track subspace drift, we adjust the core subspace every $k$ steps, defined as the subspace update interval, by minimizing its Euclidean distance to the current gradient. The objective is formalized in the cost function (2).
@@ 2 SubTrack++
where $A$ is the solution to the least squares problem. The derivative of (2) with respect to $S_{t}$ is given in (3), and the residual $R={G_{t}}-S_{t}A$ lies in the orthogonal complement of $S_{t}$. To update the subspace, we calculate the tangent vector $\nabla F$ on the Grassmannian, as shown in (4) based on Edelman et al. (1998), where the second equality holds because $R$ is orthogonal to $S_{t}S_{t}^{\top}$.
@@
The tangent vector $\nabla F$ provides the direction for adjusting the subspace by accounting for the error lying in the orthogonal complement. However, to minimize changes to the subspace, SubTrack++ first computes a rank-$1$ approximation of $\nabla F$, determined by its largest singular value and the corresponding singular vector obtained from its SVD, and represented as $\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ in the final equality of (4). This approximation is then used to update the subspace.
@@
This update rule preserves the orthonormality of $S_{t+1}$, ensuring it remains on the Grassmannian. The last term in (4), projects the previous subspace onto the orthogonal complement of $\widehat{V}_{F}$, ensuring that the portion of $S_{t}$ which has not been updated in this step is still included.
Algorithm 1 SubTrack++
(Subspace Tracking, Projection-Aware Optimizer, Recovery Scaling, Regular Adam)
[…]
  $S_{0}\leftarrow U[:,:r]$ , where $U,S,V\leftarrow\text{SVD}(G_{0})$Initializing First Subspace}
  for $t=0,\ldots,T$ do
     if $t$ mod $k==0$ then
        $G_{lr}=\arg\min_{A}\|(S_{t-1}A-G_{t})\|^{2}$, and $R=G_{t}-S_{t-1}G_{lr}$ $\nabla F=-2RG_{lr}^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ $S_{t}=(S_{t-1}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t-1}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
        $M_{t}\leftarrow\beta_{1}\cdot(S_{t}^{\top}S_{t-1}M_{t-1})+(1-\beta_{1})\cdot\widetilde{G}_{t}$$\widetilde{G}_{t}=S_{t}^{\top}G_{t}$: low-rank projection of $G_{t}$} $\mathcal{V}_{t}\leftarrow\beta_{2}\cdot[(1-\beta_{2}^{t-1})|(S_{t}^{\top}S_{t-1})^{2}\cdot(\mathcal{V}_{t-1}-M_{t-1}^{2})+(S_{t}^{\top}S_{t-1}\cdot M_{t-1})^{2}|]+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$
     else
        $S_{t}=S_{t-1}$
        $M_{t}\leftarrow\beta_{1}\cdot M_{t-1}+(1-\beta_{1})\cdot\widetilde{G}_{t}$ $\mathcal{V}_{t}\leftarrow\beta_{2}\cdot\mathcal{V}_{t-1}+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$
     end if
      $\widetilde{G}_{t}^{O}=M_{t}\oslash\sqrt{\mathcal{V}_{t}+\epsilon}$, $\widehat{G}_{t}=S_{t}\widetilde{G}_{t}^{O}$$\widetilde{G}_{t}^{O}$: optimizer’s output, $\widehat{G}_{t}$: projected-back gradients }
[…]
     if $\frac{\Lambda_{t}}{\Lambda_{t-1}}>\zeta$ then $\Lambda_{t}\leftarrow\frac{\Lambda_{t}}{\|\Lambda_{t}\|}\cdot\zeta\|\Lambda_{t-1}\|$
     $W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}$ $-\alpha\cdot\Lambda_{t}$
  end for
Projection-Aware Optimizer. In Adam, the first and second momentum update rule is shown in (6) and (7), respectively (reminder: $\widetilde{G}_{t}$ is the projection of gradient into low-rank subspace).
$\small M_{t}\leftarrow\beta_{1}\cdot M_{t-1}+(1-\beta_{1})\cdot\widetilde{G}_{t}$ (6)
$\small\mathcal{V}_{t}\leftarrow\beta_{2}\cdot\mathcal{V}_{t-1}+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$ (7)
Inspired by methods such as LDAdam (Robert et al., 2025), we emphasize the importance of updating optimizer states in a projection-aware manner to account for shifting subspaces; otherwise, misaligned projections can distort the optimizer’s performance. To address this, at each subspace update step, we modify Adam’s original update rules in (6) and (7), replacing them with projection-aware counterparts represented in (8) and (9), which reflect subspace changes into optimizer statistics (Robert et al., 2025). Further details regarding this projections can be found in Appendix C.
$\small M_{t}\leftarrow\beta_{1}\cdot(S_{t}^{\top}S_{t-1}M_{t-1})+(1-\beta_{1})\cdot\widetilde{G}_{t}$ (8)
$\small\mathcal{V}_{t}\leftarrow\beta_{2}\cdot[(1-\beta_{2}^{t-1})|(S_{t}^{\top}S_{t-1})^{2}\cdot(\mathcal{V}_{t-1}-M_{t-1}^{2})+(S_{t}^{\top}S_{t-1}\cdot M_{t-1})^{2}|]+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$ (9)
These projection-aware update rules enables Adam optimizer to track optimization dynamics precisely as the subspace evolves, achieving significantly better practical performance.
[…]
$\small W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}-\alpha\cdot\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ (10)
Here $\alpha$ is the learning rate and $\phi_{t}(G_{t})$ is the scaling factor computed based on the low-rank gradient representation, $\widetilde{G}_{t}$, and the optimizer’s processed output, $\widetilde{G}_{t}^{O}$ as:
$\small\phi_{t}(G_{t})_{i}=\frac{\|\widetilde{G}_{t,:,i}^{O}\|}{\|\widetilde{G}_{t,:,i}\|}$ (11)
As suggested by Fira (Chen et al., 2025), this additional term must be controlled using a gradient-clipping-like method to ensure training stability. Building on their observation, we also limit the growth rate of $\Lambda_{t}=\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ by a factor $\zeta$, and apply a correction whenever it exceeds this threshold as:
$\small\Lambda_{t}\leftarrow\frac{\Lambda_{t}}{\|\Lambda_{t}\|}\cdot\zeta\|\Lambda_{t-1}\|$ (12)
By incorporating a geometry-aware perspective into low-rank optimization and applying this approach across all components of the training pipeline, SubTrack++ demonstrates robust and stable training, and achieves state-of-the-art performance while maintaining minimal memory footprint and wall-time. The overall flow of operations in SubTrack++ is illustrated in Algorithm 1.